please help me solve this :the complement of an angle is 18 degree more than 3 times the measure of the angle. find the measure of both angles.
Answer by mananth(16949)

please help me solve this :the complement of an angle is 18 degree more than 3 times the measure of the angle. find the measure of both angles.
break up the word problem in parts and set up the equation

let the angle be x
the complement of an angle is 90-x
The compliment is 18 degree more than (3 times the measure of the angle).(3x)
complement = 18+3x
90-x= 18+3x
2x=72
/2
x=36 deg



find the measure of both angles.